# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities

### v1

@@ Abstract
In recent years, extensive research has been conducted to explore the utilization of machine learning algorithms in various direct-detected and self-coherent short-reach communication applications. These applications encompass a wide range of tasks, including bandwidth request prediction, signal quality monitoring, fault detection, traffic prediction, and digital signal processing (DSP)-based equalization. As a versatile approach, machine learning demonstrates the ability to address stochastic phenomena in optical systems networks where deterministic methods may fall short. However, when it comes to DSP equalization algorithms like feed-forward/decision feedback equalizers (FFE/DFE) and Volterra-based nonlinear equalizers, their performance improvements are often marginal, and their complexity is prohibitively high, especially in cost-sensitive short-reach communications scenarios such as passive optical networks (PONs). Time-series machine learning models offer distinct advantages over frequency domain models in specific contexts. They excel in capturing temporal dependencies, handling irregular or nonlinear patterns effectively, and accommodating variable time intervals. Within this extensive survey, we outline the application of machine learning techniques in short-reach communications, specifically emphasizing their utilization in high-bandwidth demanding PONs. Notably, we introduce a novel taxonomy for time-series methods employed in machine learning signal processing, providing a structured classification framework. Our taxonomy categorizes current time series methods into four distinct groups: traditional methods, Fourier convolution-based methods, transformer-based models, and time-series convolutional networks. Finally, we highlight prospective research directions within this rapidly evolving field and outline specific solutions to mitigate the complexity associated with hardware implementations. We aim to pave the way for more practical and efficient deployment of machine learning approaches in short-reach optical communication systems by addressing complexity concerns.
@@
machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities
@@ 1 Introduction
Short-reach optical transmission systems have gained substantial attraction owing to their remarkable attributes of high bandwidth and low latency Durisi (2016). In the evolving landscape of communication technologies, short-reach optical communication has emerged as an essential domain, driven by the increasing demand for high-speed data transfer in applications such as inter-data centers Kapoor (2012), access/local area networks and industrial automation XieY (2022). This increasing demand requires efficient, low-latency communication systems tailored to short-reach scenarios, typically up to 100 km. While long-haul, optical communication has been immersive in data transmission, but its applicability encounters challenges when adapting to the constraints of shorter distances. This is mainly due to physical and technical limitations that prevents its seamless integration into contemporary networking environments characterized by the need for rapid, energy-efficient, and cost-effective data transfer over limited distances. Passive optical networks (PONs) have emerged as a cost-efficient solution for access networks, effectively addressing challenges and meeting the requirements of modern networking scenarios. PONs optimize data transfer efficiency within constrained distances, providing a streamlined alternative to point-to-point communications.
Despite their efficiency, the capacity of PONs and short-reach optical systems, in general, is limited by intrinsic properties of optical components, as well as the fiber itself, for distortions such as chromatic dispersion (CD), nonlinearity, and polarization mode dispersion (PMD) Wu, Q. (2023). For instance, intensity-modulation and direct-detection (IMDD) performance is hindered by power fading attributed to fiber CD Ranzini (2019).
Ongoing research endeavors are dedicated to advancing optical detection schemes to overcome these limitations and facilitate high speed in both short-reach and long-haul optical communication networks D. Che (2016). For instance, the regeneration of coherent optical systems in the last decade has marked a breakthrough by transcending intensity-only modulation Li (2022). Coherent systems employ external modulators to imprint complex baseband signals to the optical field. The optical coherent receiver, equipped with phase diversity, linearly recovers signals and compensates for fiber impairments through digital signal processing (DSP) S. J. Savory (2007). Coherent technology enables the transmission of advanced modulation formats and polarization multiplexing to increase the signal bit rate significantly. Additionally, it triggers dense wavelength division multiplexing (DWDM) and super-channels, propelling long-haul optical networks into a capacity realm beyond multi-Terabit/s. C. R. S. Fludger (2008).
@@
In contrast to coherent transmission, IMDD operates by encoding information into the intensity of the optical signal, with the modulation signal being real-valued and positive Huang (2022). The implementation of IMDD eliminates the need for complex optical components and local oscillators, reducing hardware complexity. Additionally, IMDD systems are less susceptible to phase noise and polarization-related issues, making them robust and practical for scenarios where cost efficiency and simplicity are paramount. The advantages of IMDD, including its ease of implementation and resilience, position it as a compelling choice for optimizing short-distance optical communication Huang (2022). In addition, operational and safety factors can impose restrictions on the peak and average values of the modulated signal in IMDD. These constraints introduce distinct characteristics to its operational framework Kartal (2011). Various models, such as the Poisson channel, square-root Gaussian channel, and Gaussian channel with input-dependent noise, among others, exist to rapidly assess and characterize IMDD systems Tsiatmas (2013); Moser (2012); Safari (2015). Each of these models provides valuable insights into the intricate dynamics of this communication paradigm, enabling quick and efficient analysis. Recently, in short-reach communication systems, there has been a shift towards replacing the traditional DSP functionality with machine learning (ML) as the sole processing mechanism. In contrast to conventional methodologies that heavily depend on intricate analog components Fadlullah (2011), this strategy seamlessly integrates ML within the DSP framework Yi et al. (2018), providing a flexible and adaptable solution for enhancing optical communication over short distances. It has been reported in Kaur (2021) that the combination of ML with DSP techniques allows these systems to dynamically adapt and optimize signal parameters, effectively addressing impairments and variations in real time without requiring intricate hardware adjustments.
Furthermore, the integration of ML exclusively within the DSP domain brings streamlined and efficient implementation processes to short-reach communication systems. This approach represents a paradigm shift as it not only reduces costs associated with complex hardware setups but also emphasizes the effectiveness of intelligent signal processing solutions in short-reach optical transmission Simeone (2018); Yi et al. (2018); Rodrigues (2019).
@@
In this survey, we examine the significant advancements in short-reach communication research over the last decade. Firstly, we summarize several key research tasks (Section 2) and then focus on the equalization problem, introducing benchmark DSP methods (Section 3) and ML algorithms (Section 4). Subsequently, we categorize recent sequence models in the ML field (Section 5), dividing them into convolution-based models, transformer-based models, and Fourier-based neural networks. We explore each method’s advantages, disadvantages, and complexities in addressing the equalization problem. In the last section, we present an overview of the model compression field, outlining two approaches to compressing models. We view these approaches as potential solutions for tackling hardware complexity concerns.
@@
Our primary contribution lies in summarizing the past research on ML implementation in short-reach communications for various applications. Specifically, our contributions are the following:
@@
    We review existing deep learning models, providing a comprehensive understanding of their principles, characteristics, and hypothesis classes. This facilitates an in-depth exploration for researchers seeking supervised neural-network-based ML models suitable for their specific applications.
@@
    We highlight the features and complexities of these models, elucidating recent developments in the field of deep learning. This information is valuable for researchers interested in delving deeper into research and staying abreast of current advancements.
@@
    We discuss the current limitations and research gaps in the ongoing development of deep learning, addressing the challenges posed by these factors in real-world applications. Moreover, we proffered constructive insights regarding model selection and prospective trajectories.
@@
    Given the prevailing backdrop of high hardware complexity, we introduce a solution in the field of deep learning: the principles of model compression. We present existing works that employ this approach within the optical communication field, aiming to inspire more researchers to pursue research in this domain.
@@ 2 Applications in Short-Reach Systems
After systematically organizing recent literature in the past few years, we categorize machine learning-based short-reach system research into four classes based on application tasks: Bandwidth Request and Prediction, Subcarrier Allocation, Equalization and Fault Detection. We clarify the physical and mathematical aspects of their respective tasks, enumerate several recent works, and provide a summary of current advancements.
@@
Bandwidth Request and Prediction: It aims to leverage network information to predict future bandwidth availability and enable its utilization by related applications. In mathematical terms, the real-time bandwidth forecast at a specific time (t) involves estimating the available bandwidth that will be accessible in the immediate future (t + $\tau$) Qi-yuZ (2012). One proposed method, known as predictive-dynamic bandwidth allocation (P-DBA), utilizes this concept to predict high-priority traffic during waiting periods, resulting in reduced latency and packet loss rates within Gigabit PON (GPON) Qi-yuZ (2012). Another approach demonstrated in PanagiotisS (2016) leverages the k-nearest neighbor algorithm to predict additional bandwidth requirements for each optical network unit (ONU) in a PON. This adaptive learning-based approach dynamically adjusts the k value based on real-time traffic conditions, showcasing the adaptability of ML in optimizing bandwidth allocation PanagiotisS (2016). Artificial neural networks (ANNs) have also shown promise in achieving flexible bandwidth allocations across various application scenarios, particularly emphasizing low-latency objectives LihuaR (2018); YiL (2019). For example, feed-forward-based ANNs, explored in Mikaeil (2018), are utilized to predict packet arrivals in time-division multiple access (TDMA) ONUs, effectively reducing additional DBA processing delays Mikaeil (2018). Furthermore, in Ye (2018), XgboostXgboost (2016) is employed to predict bandwidth requests for ONUs in Ethernet PON (EPON), optimizing bandwidth utilization across polling periods. This study introduces a dynamic wavelength and bandwidth assignment scheme for time and WDM (TWDM) PONs, incorporating regression techniques for efficient resource allocation Ye (2018). Collectively, these studies highlight the versatility of ML approaches in addressing bandwidth prediction and request challenges, paving the way for more adaptive and efficient short-reach optical communication systems Qi-yuZ (2012); PanagiotisS (2016); LihuaR (2018); YiL (2019); Mikaeil (2018); Ye (2018).
@@
Subcarrier Allocation: The optimization of bandwidth allocation for enhanced spectral efficiency has led to increased interest in subcarrier allocation for PONs. This approach involves mathematically formulating the allocation problem as an integer linear programming (ILP) task, which includes tasks such as optimizing wavelength configurations, assigning subcarriers to transmitters, and minimizing lost traffic and energy costs. To address this challenge, deep reinforcement learning has emerged as a promising technique that enables dynamic subcarrier sharing among ONUs, facilitating efficient DBA. At the medium access control (MAC) layer, the dynamic subcarrier allocation (DSA) algorithm schedules ONU upstream transmissions by considering instantaneous bandwidth requirements and existing traffic conditions KonstantinosK (2012) . This showcases the adaptability of ML in resource scheduling. Several studies focus on algorithm-level cost reduction and two-dimensional resource scheduling for orthogonal frequency-division multiplexing (OFDM)-PONs including KonstantinosK (2012); WansuL (2017); MeihuaB (2013). These DSA algorithms address challenges related to latency, throughput, and energy efficiency, highlighting the versatility of ML in enhancing subcarrier allocation strategies ZhuM (2022). Moreover, the integration of traffic prediction technology and fair-aware DSA algorithms, as proposed in YumikoS (2017) and ZhuM (2022), further enhances the performance of subcarrier allocation in short-reach optical communication systems. These advancements improve the efficiency and adaptability of subcarrier allocation by applying ML methodologies YNakayama (2022).
Power Budget Problem: The electric power budgeting problem is focusing on predicting energy consumption for future time intervals with the given historical data on power consumption and relevant environmental factors such as weather conditions, user behavior, and equipment efficiency, the goal is to predict for future time intervals. However, the large-scale systematic ML development is largely constrained by the limited number of publicly available datasets, and recent work has provided a basic data construction process and baseline model Alberto C (2023).
@@
Equalization: The objective of this task is to minimize fiber-induced distortions by employing post-processing techniques that compensate for linear effects, such as chromatic dispersion (CD). Mathematically, the equalizer optimizes the function $f(x)$ to ensure that the equalized output sequence $y$ closely approximates the input signal. Performance evaluation primarily relies on the bit error ratio (BER). In PON systems, the utilization of post-equalizers that employ shallow-based deep learning models shows promise in mitigating nonlinear distortions, especially in scenarios involving modulator nonlinearities or high launched optical power to meet stringent power budgets Li (2022). In the context of PONs, these situations can give rise to self-phase modulation (SPM) in single-channel scenarios or, in the case of multi-channel PONs, additional phenomena such as cross-phase modulation (XPM) and four-wave mixing (FWM). In IMDD PON, low-complexity ANN-based equalizers have demonstrated comparable performance to Volterra-based equalizers in PAM4 systems ZhiweiChen (2022). While post-equalization techniques have proven effective, the computational complexity at the receiver side poses challenges for the ONU. To address this issue, strategies for centralized pre-equalization at the transmitter side have been proposed. Examples of these strategies include memory polynomial-based pre-equalizers ZhiweiChen (2022) and trained NN-based pre-equalizers XueL (2021). These methods enhance equalization effectiveness while maintaining the simplicity of the DSP at the ONU receiver. These advancements highlight the potential of ML in optimizing equalization strategies in PONs and other short-reach systems. This survey provides a comprehensive overview of existing equalization techniques,focusing on deep learning methods that have demonstrated effectiveness but have not yet been widely adopted.
@@
Fault Detection: Short-reach optical communication systems, including PONs, are susceptible to failures such as fiber cuts, equipment failures, power outages, and natural disasters and ONU transceiver malfunctions Abdelli (2023). The impact of service disruptions can result in significant financial losses for service providers. Identifying faulty ONUs presents challenges, especially in scenarios where nearly equidistant branch terminations lead to overlapping reflections, making it difficult to pinpoint the exact defective branch Abdelli (2023). Conventional monitoring approaches become less reliable as PON systems grow in complexity. Recent advancements in ML-enabled proactive fault monitoring offer promising solutions to ensure stable network operation. ML-based fault prediction algorithms utilize past network fault data to discover underlying patterns and similarities among events. By doing so, these algorithms enhance the detection of optical network problems and facilitate proactive repairs, thereby avoiding potential issues from occurring.. Several research papers propose using ML algorithms for monitoring management in optical networks. Notably, technologies like random forest and ANN algorithms have been employed to continuously monitor the BER, predict network component failures, and assess fault severity APVela (2018). Wang et al. introduced a hybrid approach combining double exponential smoothing and support vector machines for equipment failure prediction in software-defined metropolitan area networks Wang (201). Bayesian network-based models have also been developed for diagnosing PON faults Wang (201).

### v2

# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error ratio, optical signal-to-noise ratio, nonlinearities

@@ Abstract
Recently, extensive research has been conducted to explore the utilization of machine learning (ML) algorithms in various direct-detected and (self)-coherent short-reach communication applications. These applications encompass a wide range of tasks, including bandwidth request prediction, signal quality monitoring, fault detection, traffic prediction, and digital signal processing (DSP)-based equalization. As a versatile approach, ML demonstrates the ability to address stochastic phenomena in optical systems networks where deterministic methods may fall short. However, when it comes to DSP equalization algorithms such as feed-forward/decision feedback equalizers (FFE/DFE) and Volterra-based nonlinear equalizers, their performance improvements are often marginal, and their complexity is prohibitively high, especially in cost-sensitive short-reach communications scenarios such as passive optical networks (PONs). Time-series ML models offer distinct advantages over frequency domain models in specific contexts. They excel in capturing temporal dependencies, handling irregular or nonlinear patterns effectively, and accommodating variable time intervals. Within this survey, we outline the application of ML techniques in short-reach communications, specifically emphasizing their utilization in high-bandwidth demanding PONs. We introduce a novel taxonomy for time-series methods employed in ML signal processing, providing a structured classification framework. Our taxonomy categorizes current time series methods into four distinct groups: traditional methods, Fourier convolution-based methods, transformer-based models, and time-series convolutional networks. Finally, we highlight prospective research directions within this rapidly evolving field and outline specific solutions to mitigate the complexity associated with hardware implementations. We aim to pave the way for more practical and efficient deployment of ML approaches in short-reach optical communication systems by addressing complexity concerns.
@@
machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error ratio, optical signal-to-noise ratio, nonlinearities
@@ 1 Introduction
Short-reach optical transmission systems have gained substantial attraction owing to their remarkable attributes of high bandwidth and low latency Durisi (2016). In the evolving landscape of communication technologies, short-reach optical communication has emerged as an essential domain, driven by the increasing demand for high-speed data transfer in applications such as inter-data centers Kapoor (2012), access/local area networks and industrial automation XieY (2022). This increasing demand requires efficient, low-latency communication systems tailored to short-reach scenarios, typically up to 100 km. While long-haul, optical communication has been immersive in data transmission, its applicability encounters challenges when adapting to the constraints of shorter distances. This is mainly due to physical and technical limitations that prevent its seamless integration into existing networking environments characterized by the need for energy-efficient, and cost-effective data transmission over limited distances. Passive optical networks (PONs) utilize passive optical splitters and combiners, which are less expensive than the active components required in traditional point-to-point fiber networks. This makes PONs a cost-effective fiber-optic solution.
Since PONs rely on passive optical splitters, they inherently introduce power losses, limiting the overall power budget and the number of users that can be supported on a single PON. In addition, effects caused by the fiber, such as chromatic dispersion (CD) and nonlinearity can limit the PON-reach Wu, Q. (2023), especially when intensity-modulated and direct-detected (IMDD) high baud-rate signals are considered Ranzini (2019).
Ongoing research endeavors are dedicated to advancing optical detection schemes to overcome these limitations and increase the signal bit-rate in both short-reach and long-haul optical communication networks D. Che (2016). For instance, the regeneration of coherent optical systems in the last decade has been a major breakthrough, as they have gone beyond just using intensity-only modulation Li (2022). Coherent systems employ external modulators to employ complex baseband signals to the optical field. The optical coherent receiver, equipped with phase diversity, linearly recovers signals and compensates for fiber impairments through digital signal processing (DSP) S. J. Savory (2007). Coherent technology enables the transmission of advanced modulation formats and polarization multiplexing to increase the signal bit-rate significantly. Additionally, coherent optical systems enable dense wavelength division multiplexing (DWDM) and super-channels, which are pushing long-distance optical networks into the multi-terabit per second capacity range C. R. S. Fludger (2008).
@@
In contrast to coherent transmission, IMDD operates by encoding information into the intensity of the optical signal, with the modulation signal being real-valued and positive Huang (2022). The implementation of IMDD eliminates the need for complex optical components and local oscillators, reducing hardware complexity. Additionally, IMDD systems are less susceptible to phase noise and polarization-related issues, making them robust and practical for scenarios where cost efficiency and simplicity are paramount Huang (2022). Furthermore, practical considerations like operation and safety can limit the highest and average values of the modulated signal in IMDD systems. These restrictions give IMDD systems specific characteristics in how they function Kartal (2011). Various models, such as the Poisson channel, square-root Gaussian channel, and Gaussian channel with input-dependent noise, among others, exist to rapidly assess and characterize IMDD systems Tsiatmas (2013); Moser (2012); Safari (2015). In contrast to conventional methodologies that depend on analog components and processing Fadlullah (2011), IMDD can potentially integrate machine learning (ML) algorithms at the receiver DSP if required Yi et al. (2018), providing a flexible and adaptable solution for enhancing the transmission performance. According to Kaur (2021), the combination of ML and DSP techniques allows IMDD systems to dynamically adapt and optimize signal parameters. This addresses impairments and variations in real-time without needing complex hardware adjustments. This approach represents a significant benefit, as it not only reduces the costs associated with complex hardware setups in short-reach systems, but also highlights the effectiveness of intelligent signal processing Simeone (2018); Yi et al. (2018); Rodrigues (2019).
@@
In this survey, we examine the significant progress made in short-distance optical communications research over the past decade. First, we summarize several key research areas (Section 2). Afterwards, we focus on the equalization problem, introducing benchmark DSP methods (Section 3) and ML algorithms (Section 4). Then, we categorize recent sequence models in the ML field (Section 5), dividing them into convolution-based, transformer-based, and Fourier-based neural networks. We explore the advantages, disadvantages, and complexities of each method in addressing the equalization problem. In the final section, we provide an overview of the model compression field, outlining two approaches to compress models. We see these approaches as potential solutions for addressing hardware complexity concerns.
@@
The primary contribution of this survey is to summarize the existing research on ML implementations for short-reach optical communications across a range of applications. Specifically, our contributions are the following:
@@
    We review existing Deep Learning (DL) models, providing a comprehensive understanding of their principles, characteristics, and hypothesis classes. This facilitates an in-depth exploration for researchers seeking supervised neural-network-based ML models suitable for their specific applications.
@@
    We highlight the features and complexities of these models, elucidating recent developments in the field of DL. This information is valuable for researchers interested in delving deeper into research and staying abreast of current advancements.
@@
    We discuss the current limitations and research gaps in the ongoing development of DL, addressing the challenges posed by these factors in real-world applications. Furthermore, we provide constructive insights regarding the selection of models and potential future directions.
@@
    Given the challenge of high hardware complexity, we introduce model compression as a potential solution from the DL field. We present existing works that employ this approach within the optical communication field, aiming to inspire more researchers to pursue research in this domain.
@@ 2 Applications in Short-Reach Systems
After systematically organizing recent literature in the past few years, we have categorized ML-based research for short-reach optical systems into four classes based on application tasks: Bandwidth Request and Prediction, Subcarrier Allocation, Equalization, and Fault Detection. We clarify the physical and mathematical aspects of their respective tasks, enumerate several recent works, and provide a summary of current advancements.
@@
Bandwidth Request and Prediction: It aims to leverage network information to predict future bandwidth availability and enable its utilization by related applications. In mathematical terms, the real-time bandwidth forecast at a specific time (t) involves estimating the available bandwidth that will be accessible in the immediate future (t + $\tau$) Qi-yuZ (2012). One proposed method, known as predictive-dynamic bandwidth allocation (P-DBA), utilizes this concept to predict high-priority traffic during waiting periods, resulting in reduced latency and packet loss rates within a Gigabit PON (GPON) Qi-yuZ (2012). Another approach demonstrated in PanagiotisS (2016) leverages the k-nearest neighbor algorithm to predict additional bandwidth requirements for each optical network unit (ONU) in a PON. This adaptive learning-based approach dynamically adjusts the k value based on real-time traffic conditions, showcasing the adaptability of ML in optimizing bandwidth allocation PanagiotisS (2016). Artificial neural networks (ANNs) have also shown promise in achieving flexible bandwidth allocations across various application scenarios, particularly emphasizing low-latency objectives LihuaR (2018); YiL (2019). For example, feed-forward-based ANNs, explored in Mikaeil (2018), are utilized to predict packet arrivals in time-division multiple access (TDMA) ONUs, effectively reducing additional DBA processing delays Mikaeil (2018). Furthermore, XgboostXgboost (2016) is employed to predict bandwidth requests for ONUs in Ethernet PON (EPON), optimizing bandwidth utilization across polling periods. This study introduced a dynamic wavelength and bandwidth assignment scheme for time and WDM (TWDM) PONs, incorporating regression techniques for efficient resource allocation Ye (2018). Recent studies show that ML approaches are versatile in addressing challenges related to predicting and managing bandwidth needs. This paves the way for developing more adaptive and efficient short-reach optical communication systems in the near future Qi-yuZ (2012); PanagiotisS (2016); LihuaR (2018); YiL (2019); Mikaeil (2018); Ye (2018).
@@
Subcarrier Allocation: The optimization of bandwidth allocation for enhanced spectral efficiency has led to increased interest in subcarrier allocation for PONs. This approach involves mathematically formulating the allocation problem as an integer linear programming (ILP) task, which includes tasks such as optimizing wavelength configurations, assigning subcarriers to transmitters, and minimizing lost traffic and energy costs. To address this challenge, deep reinforcement learning has emerged as a promising technique that enables dynamic subcarrier sharing among ONUs, facilitating efficient DBA. At the medium access control (MAC) layer, the dynamic subcarrier allocation (DSA) algorithm schedules ONU upstream transmissions by considering instantaneous bandwidth requirements and existing traffic conditions KonstantinosK (2012) . This showcases the adaptability of ML in resource scheduling. Several studies focus on algorithm-level cost reduction and two-dimensional resource scheduling for orthogonal frequency-division multiplexing (OFDM)-PONs including KonstantinosK (2012); WansuL (2017); MeihuaB (2013). These DSA algorithms address challenges related to latency, throughput, and energy efficiency, highlighting the versatility of ML in enhancing subcarrier allocation strategies ZhuM (2022). Moreover, the integration of traffic prediction technology and fair-aware DSA algorithms, as proposed in ZhuM (2022) and YumikoS (2017), further enhances the performance of subcarrier allocation in short-reach optical communication systems. These advancements improve the efficiency and adaptability of subcarrier allocation by applying ML methodologies YNakayama (2022).
Power Budget Limitations: The electric power budgeting issue is about predicting future energy consumption using historical data on power usage and related environmental factors like weather, user behavior, and equipment efficiency. The goal is to forecast power consumption for upcoming time periods. However, the development of large-scale, systematic ML models for this task is limited by the lack of publicly available datasets. Recent research has provided a basic process for constructing the necessary data and has also presented baseline ML models as a starting point. Specifically, the data construction process involves compiling and organizing relevant datasets, including time-series power consumption data, weather information, occupancy patterns, and equipment performance metrics. This standardized data can then be used to develop and test ML models for power consumption forecasting. For instance, the recent work in Alberto C (2023) has introduced baseline ML models that demonstrate the feasibility of using these techniques to predict future power consumption, despite the constraints posed by the scarcity of publicly accessible datasets.
@@
Equalization: The objective of this task is to minimize fiber-induced distortions by employing post-processing techniques that compensate for linear effects, such as CD. Mathematically, the equalizer optimizes the function $f(x)$ to ensure that the equalized output sequence $y$ closely approximates the input signal. Performance evaluation primarily relies on the bit error ratio (BER). In PON systems, using shallow-based DL models for post-equalizers has shown potential in addressing nonlinear distortions for both IMDD and coherent signals. This is especially useful in scenarios with modulator nonlinearities or high-launched optical power to meet tight power budgets Li (2022). As the fiber-induced nonlinear effects are increasing in the latter case, in single-channel coherent PONs, this results in self-phase modulation (SPM). In multi-channel PONs, the increased nonlinear effects result in cross-phase modulation (XPM) and four-wave mixing (FWM). In IMDD PONs, low-complexity artificial neural network (ANN)-based equalizers have demonstrated performance comparable to Volterra-based equalizers in pulse amplitude modulation with 4 levels (PAM4) systems ZhiweiChen (2022). While post-equalization techniques have proven effective, the computational complexity at the ONU receiver is a challenge. To address this, strategies for centralized pre-equalization at the transmitter side have been proposed. Examples include memory polynomial-based pre-equalizers ZhiweiChen (2022) and trained neural network-based pre-equalizers XueL (2021). These methods enhance equalization effectiveness while keeping the ONU receiver simple.
@@
Fault Detection: Short-reach optical communication systems, including PONs, are susceptible to failures such as fiber cuts, equipment failures, power outages, natural disasters, and ONU transceiver malfunctions Abdelli (2023). Service disruptions can result in significant financial losses for service providers. Identifying faulty ONUs presents challenges, especially when nearly equidistant branch terminations lead to overlapping reflections, making it difficult to pinpoint the exact defective branch Abdelli (2023). Conventional monitoring approaches become less reliable as PON systems grow in complexity. Recent advancements in ML-enabled proactive fault monitoring offer promising solutions to ensure stable network operation. ML-based fault prediction algorithms utilize past network fault data to discover underlying patterns and similarities. By doing so, these algorithms enhance the detection of optical network problems and facilitate proactive repairs, thereby preventing potential issues from occurring. Several research papers propose using ML algorithms for monitoring management in optical networks. Notably, technologies like random forest and ANN algorithms have been employed to continuously monitor the BER, predict network component failures, and assess fault severity APVela (2018). Wang et al. Wang (201) introduced a hybrid approach combining double exponential smoothing and support vector machines for equipment failure prediction in software-defined metropolitan area networks. Bayesian network-based models have also been developed for diagnosing PON faults Wang (201).